OpenStudy (anonymous):

Find the difference. 8/x^2+9x (minus) 3/x^2-81

OpenStudy (callisto):

\[\frac{8}{x^2+9x}-\frac{ 3}{x^2-81}\]Factor the denominators \[=\frac{8}{x(x+9)}-\frac{ 3}{(x-9)(x+9)}\]\[=\frac{8(x-9)}{x(x+9)(x-9)}-\frac{ 3x}{x(x-9)(x+9)}\]\[=\frac{8x-72- 3x}{x(x-9)(x+9)}\]Can you simplify it?

OpenStudy (anonymous):

I can't :/ I'm horrible at it, and this test is timed. Please do it for me. :P

OpenStudy (callisto):

We're not supposed to help you cheat... At least you should try?

OpenStudy (anonymous):

I've been trying for about an hour. I NEED help

OpenStudy (callisto):

Not this question... I believe... You've been asking a number of questions of the same type... Here's the hint: Can you simplify the numerator? Numerator: 8x - 72 -3x

OpenStudy (anonymous):

I don't have time for this. I'm sorry.